OpenStudy (hothotleg):

I need these answered and explained, please! I have exactly 7 questions! I will fan and medal!

OpenStudy (hothotleg):

@haunting_beauty @Howard-Wolowitz

OpenStudy (anonymous):

Please post your questions one at a time.

OpenStudy (hothotleg):

Alright! So should I post the first one right now?

OpenStudy (anonymous):

yes

OpenStudy (hothotleg):

Alright 1. For the system below, solve using linear combinations method. Show all work. System: 2a+3b=1 −2a+b=11 Do you understand? Because I unfortunately don't.. D:

OpenStudy (anonymous):

@youngstudier

OpenStudy (anonymous):

@Compassionate

OpenStudy (anonymous):

@AlexandervonHumboldt2

OpenStudy (anonymous):

What is this linear combinations method?

OpenStudy (anonymous):

ok we must get a and b by themself for both equations

OpenStudy (hothotleg):

Alright, how do we do that?

OpenStudy (anonymous):

2a+3b(-3b)=1(-3b) this makes it 2a=1-3b then for your other equation we do the same −2a+b=11 −2a+b(-b)=11(-b) this makes it -2a=11-b

OpenStudy (anonymous):

ill show you the next step once you understand that

OpenStudy (hothotleg):

Alright, Let me read, I'll tell you when I am done, thank you! ^u^

OpenStudy (hothotleg):

Alright! I am done reading, I am ready for the next step!

OpenStudy (hothotleg):

:)

OpenStudy (anonymous):

ok I made a small mistake on part one of this step for -2a+b=11 you do it like this -2a(+2a)+b=11(+2a) so we have b=11+2a for the other eqation we now have 2a=1-3b we want a by its self so 2a/a=(1-3b)/2 a=1/2-3b/2 let me know when your ready for the next step :)

OpenStudy (hothotleg):

Alright! That was surprisingly easy! Thank you! I am ready for the next step.

OpenStudy (hothotleg):

:)

OpenStudy (anonymous):

ok now plug the equations together b=11+2(1/2-3/2(b)) using the distributive property multiply 2 and whats in the () b=11+2(1/2)-3/2(b)(2)) b=11+1-3b b=12-3b now get b on one side of the = b(+3b)=12-3b(+3b) 4b=12 4b/4=12/4 b=3

OpenStudy (anonymous):

@hothotlEG let me know when your ready for the last step this is easy :)

OpenStudy (hothotleg):

:) thank you so much! this is actually fun! I'm ready for the next step!

OpenStudy (hothotleg):

:)

OpenStudy (anonymous):

ok since we have b=3 we can plug it into either equation lets use 2a+3b=1 2a+3(3)=1 2a+9=1 2a+9(-9)=1(-9) 2a=-8 2a/2=-8/2 a=-4 and if you want to check my math plug in the a=-4 and b=3 in to both of your equations and they should be the same answer. :)

OpenStudy (anonymous):

@hothotlEG

OpenStudy (hothotleg):

Thank you very much! Your very good at this!

OpenStudy (hothotleg):

I need to go now, but again, thank you so much!

OpenStudy (anonymous):

np